# Follow the Mass - A Concordance Picture of Tidal Disruption Events

### v1

@@ Abstract
Three recent global simulations of tidal disruption events (TDEs) have produced, using different numerical techniques and parameters, very similar pictures of their dynamics. In typical TDEs, after the star is disrupted by a supermassive black hole, the bound portion of the stellar debris follows highly eccentric trajectories, reaching apocenters of several thousand gravitational radii. Only a very small fraction is captured upon returning to the vicinity of the supermassive black hole. Nearly all the debris returns to the apocenter, where shocks produce a thick irregular cloud on this radial scale and power the optical/UV flare. These simulation results imply that over a few years, the thick cloud settles into an accretion flow responsible for the long term emission. Despite not being designed to match observations, the dynamical picture given by the three simulations aligns well with observations of typical events, correctly predicting the flares’ total radiated energy, luminosity, temperature and emission line width. On the basis of these predictions, we provide an updated method (TDEmass) to infer the stellar and black hole masses from a flare’s peak luminosity and temperature. This picture also correctly predicts the luminosity observed years after the flare. In addition, we show that in a magnitude-limited survey, if the intrinsic rate of TDEs is independent of black hole mass, the detected events will preferentially have black hole masses $\sim 10^{6\pm 0.3}M_{\odot}$ and stellar masses of $\sim 1-1.5M_{\odot}$.
[…]
patch/longtable=false
@@ 1 Introduction
Tidal disruption events (TDEs) are inherently dramatic: a star is ripped apart by the gravity of a supermassive black hole. The result is a flare in which, for a month or two, the former star radiates with $\sim 10^{10}\times$ its ordinary luminosity. Over the duration of the flare (typically months), TDEs are among the most luminous transients known. Once hard to find, there are now $\sim 100$ known examples (Blanton et al., 2017; Shappee et al., 2014; Kaiser et al., 2002; Bellm et al., 2019; Hammerstein et al., 2023), and many more can be expected from new instruments (e.g., the Rubin Observatory: Bricman & Gomboc 2020; Szekerczes et al. 2024 and ULTRASAT: Ben-Ami et al. 2022) due to go online soon.
@@
To identify where and how the radiation is emitted, the first task is to track where the debris mass goes. When the star’s orbit is effectively parabolic and the star has been fully disrupted, half the star’s mass is unbound and escapes to infinity, while the other half remains bound to the black hole (Rees, 1988). Although there is a firm consensus that the bound mass is initially placed on highly-eccentric orbits whose apocenters are thousands of gravitational radii, $\sim O(10^{2})\times$ the stellar pericenter, hitherto there has been little agreement about its whereabouts following its first return to the pericenter. The oldest, and still most prevalent, view has been that the bound material immediately forms a compact accretion disk on the pericenter scale. This picture predicts a very luminous X-ray source, but the observed spectra have luminosities $\sim 10^{-2}\times$ this prediction, and shapes well fit by black bodies with temperatures $\gtrsim 10^{4}$ K and radiating areas corresponding to the apocenter scale (see van Velzen et al. (2020); Hammerstein et al. (2023) for two recent data compilations). To explain these unexpected observations, many (e.g., Strubbe & Quataert (2009, 2011); Metzger & Stone (2016)) have suggested that the X-rays are reprocessed by different material associated with the event whose photosphere happens to be at this distance from the black hole. Alternatively, it has also been suggested (Shiokawa et al., 2015) that, rather than forming a compact disk, the bound material returns to the apocenter region and, upon colliding with later-returning bound debris, forms a large irregular cloud on a scale similar to the apocenter distance. In this picture (Piran et al., 2015), the shocks associated with these collisions power the flare.
@@
Recently, the numerical barriers have been overcome by three different groups (Ryu et al., 2023b; Steinberg & Stone, 2024; Price et al., 2024) using very different numerical methods. In this paper, we will show that these simulations largely agree on where the bound debris goes after its first return to the vicinity of the black hole. In addition, we will outline the striking observational implications stemming from this consensus. These implications touch upon the total radiated energy during the flare peak and its spectral shape, the long-term evolution of TDE emission, and how the event rates in flux-limited samples depend on the stellar and black hole masses. We will first briefly summarize the relevant observational results, focusing on TDEs of the most commonly seen variety, those dominated by an optical/UV flare (Sec. 2). Next (in Sec. 3) we will remind readers of the tidal debris’ characteristic lengthscales, orbital energies, etc. Following these presentations of background, we will summarize the three new global simulations (Sec. 4). The heart of our work will appear in Sec. 5, where we demonstrate how the shared results of the simulations lead to important statements about both the underlying dynamics of TDEs and many of their observed properties. Our conclusions will then be summarized in Sec. 6.
@@ 2.1 The optical flare
In most optical/UV TDEs, the optical/UV flux rises quickly (over a few weeks) to a peak of $\sim 3\times 10^{43}-3\times 10^{44}$ erg s-1, followed by a lengthy (a few months) decline. The spectra are well-fit by single-temperature Planck functions, whose typical temperatures are $\sim 3\times 10^{4}$ K and corresponding black-body emitting radii are $10^{14-15}$ cm (van Velzen et al., 2020; Hammerstein et al., 2023). When lines are observed, their typical width corresponds to $\sim 5000$ km/sec (see e.g., Gezari, 2021, for a review). For a black hole mass $\sim 10^{6}M_{\odot}$, the scale of the emitting area is consistent with the scale of orbits with the speed inferred from the lines.
In some cases, the luminosity declines from its peak $\propto t^{-\alpha}$ with $\alpha\simeq 5/3$, the expected slope of the mass infall rate’s decline (Rees, 1988; Phinney, 1989). However, the entire sample exhibits a wider range of lightcurve power-law indices, $1\lesssim\alpha\lesssim 3$ (Hammerstein et al., 2023). Combining lightcurve and single-observation luminosity data, one finds the radiated energy is $\sim 10^{50.5\pm 0.5}$ erg. At late times, years or more after the peak, the decay is much shallower. For those events observed $\gtrsim 0.5$ yr past the peak, although there is a good deal of scatter, the bolometric luminosity typically falls below the peak by a factor $\sim 10$ (van Velzen et al., 2019a, 2021). The late-time radiated energy is therefore comparable to the prompt radiated energy.
@@ 2.2 X-rays
Although there are a few dozen examples of TDEs whose peak X-ray luminosities are comparable to the peak optical/UV luminosities often seen, only a few TDEs discovered by optical/UV flaring have been associated with X-ray flares (Saxton et al., 2021). Nonetheless, the great majority of TDEs seen to produce X-rays, whether optically-detected or not, have rather soft spectra during the flare: when fitted by a Planck function, the characteristic temperature $kT\sim 50-100$ eV (Saxton et al., 2021). There is, however, a good deal of diversity in their lightcurves: in terms of how rapidly the X-rays decline (Auchettl et al., 2017), non-monotonic behavior, and separation in time from any associated optical/UV flare (Jonker et al., 2020; Malyali et al., 2024; Wevers et al., 2021, 2024) or spectral changes (Jonker et al., 2020; Sazonov et al., 2021; Khorunzhev et al., 2022).
@@ 3.1 Lengthscales
where $m_{\star}$ is the stellar mass in solar masses, $m_{\rm BH,6}$ is the black hole mass in units of $10^{6}M_{\odot}$, and $r_{\rm g}\equiv GM_{\rm BH}/c^{2}$ is the black hole’s gravitational radius. Here and in the rest of the text, we describe the main-sequence mass-radius relation by the power-law $R_{*}=0.93R_{\odot}m_{\star}^{0.88}$ (Ryu et al., 2020c).
whenever $m_{\star}$ appears in a scaling relation, if its exponent is written as a decimal quantity, part of the $m_{\star}$-dependence is through $R_{*}(m_{\star})$.
The tidal radius estimator $r_{\rm t}$ is based on an order-of-magnitude argument. Other radii relevant to tidal disruptions differ from it by factors of order unity. In particular, the critical radius within which a star can be completely disrupted is ${\cal R}_{T}=\Psi(M_{\star},M_{{\rm BH}})r_{\rm t}$ (Ryu et al., 2020c, d). The correction factor $\Psi(M_{\star},M_{{\rm BH}})$ is defined in the Appendix. Although $\Psi(M_{\star},M_{{\rm BH}})$ is of order unity, its dependence on the stellar mass and the black hole mass is important when estimating TDE rates.
@@ 3.2 Energy scales and orbital properties
In other words, the orbital energy of the debris is, for typical parameters, $\sim 10^{-2}$ the potential energy near the nominal tidal radius. However, once again, detailed calculations (Ryu et al., 2020c, d; Law-Smith et al., 2020) have found that consideration of the internal density profile of main sequence stars introduces order-unity correction factors: Ryu et al. (2020c) defined them by $\Xi=\Delta E/\Delta E_{0}$, where the energy range $-\Delta E<E<+\Delta E$ contains 90% of the debris mass. Details about the correction factor $\Xi$, including fitting formulæ for its dependence on $M_{\star}$ and $M_{{}_{\rm BH}}$, are provided in the Appendix  A.
@@
As $m_{\star}$ increases, the internal density profile of main sequence stars becomes increasingly centrally-concentrated. This causes $\Xi$ to rise as a function of $M_{\star}$, and the net result is for there to be almost no net trend in $t_{0}$ as $M_{\star}$ increases from $\simeq 0.1M_{\odot}$ to $\simeq 10M_{\odot}$ (Ryu et al., 2020c). On the other hand, the explicit scaling with $M_{\rm BH}$ is augmented by the implicit dependence on $M_{\rm BH}$ through $\Xi$, making $t_{0}\propto m_{{}_{\rm BH,6}}^{0.6}$.
@@ 3.3 Circularization and the inverse energy crisis
where $\sigma_{\rm SB}$ is the Stefan-Boltzmann constant. This is 50$\sim$ 50$\times$ 50 × the observed optical/UV temperature (van Velzen et al., 2020; Hammerstein et al., 2023). Typical observed velocities from motion at a few tens of $r_{\rm g}$ would be $\sim 30,000$ km s-1, about ten times larger than implied by the measured emission line widths.
@@ 4.1 Realistic, Global Simulations
Three simulations stand out as satisfying all these criteria. We list them in order of their publication dates. Ryu et al. (2023b) employed two fixed grids, one a small box following the star’s orbit, the other a large volume surrounding the black hole. Simultaneous evolution on the two grids was coordinated through a “multipatch" system (Shiokawa et al., 2018; Avara et al., 2023). The programs running on both grids solve the general relativistic hydrodynamics equations; the one responsible for the star adds relativistically-consistent stellar self-gravity. Ryu et al. (2023b) considered a $3M_{\odot}$ star, whose internal structure was taken from a MESA model of a middle-aged main-sequence star, and a $10^{5}M_{\odot}$ black hole. The equation of state included LTE radiation pressure, but there was no computation of radiation transfer because, for these parameters, the cooling time was always very long compared to the evolution time. The duration of this simulation was $3t_{0}$. Steinberg & Stone (2024) solved the equations of Newtonian hydrodynamics and radiation transport in the flux-limited diffusion approximation on a moving mesh. Stellar self-gravity was computed via a quadrupole moment tree, and the black hole’s gravity was approximated as a Paczynski-Wiita potential with a softening term. They chose a $1M_{\odot}$ star and a $10^{6}M_{\odot}$ black hole, but described the internal structure of the star as an $n=3/2$ polytrope, i.e., an isentropic structure for a gas with an adiabatic index of 5/3. This simulation ran for $\approx 1.4t_{0}$. The third (Price et al., 2024) was produced by an SPH code with relativistic hydrodynamics in a Schwarzschild spacetime and the same equation of state as in Ryu et al. (2023b). Their star’s initial structure was the same as in Steinberg & Stone (2024), but its self-gravity was calculated as an integral over the Newtonian Green’s function for the Poisson Equation with a softening term to smooth short lengthscale fluctuations. The simulation covered a time up to $\simeq 9t_{0}$.
@@ 4.2 Consensus dynamical results
Debris that has passed through the nozzle shock swings back out toward an apocenter that is slightly smaller because of the orbital energy dissipated in the shock; its natal specific angular momentum is so small that the angular momentum exchange in the shock hardly affects the apocenter. Because the disruption of the star takes place across a range of distances from the black hole (roughly from $r\simeq r_{\rm p}$ to $r\simeq 20r_{\rm p}$: Ryu et al. (2020d)) and because of relativistic effects in both the tidal stress and the orbits, the lines of apsides of the debris orbits stretch across a range of angles $\sim 10^{\circ}$ (Shiokawa et al., 2015). When the range of angles is this small and the orbits are highly eccentric, the streams intersect near apocenter (see also the related geometric argument of Dai et al. (2015)). Two views of these shocks are presented in Figure 1, which highlights shock locations by portraying temperature (see also Fig. 1 of Steinberg & Stone (2024)).

### v2

@@ Abstract
Three recent global simulations of tidal disruption events (TDEs) have produced, using different numerical techniques and parameters, very similar pictures of their dynamics. In typical TDEs, after the star is disrupted by a supermassive black hole, the bound portion of the stellar debris follows highly eccentric trajectories, reaching apocenters of several thousand gravitational radii. Only a very small fraction is captured upon returning to the vicinity of the supermassive black hole. Nearly all the debris returns to the apocenter, where shocks produce a thick irregular cloud on this radial scale and power the optical/UV flare. These simulation results imply that over a few years, the thick cloud settles into an accretion flow responsible for the long term emission. Despite not being designed to match observations, and without adjusting any parameters, the dynamical picture given by the three simulations aligns well with observations of typical events, correctly predicting the flares’ typical total radiated energy, luminosity, temperature and emission line width. On the basis of these predictions, we provide an updated method (TDEmass) to infer the stellar and black hole masses from a flare’s peak luminosity and temperature. This picture also correctly predicts that the luminosity observed years after the flare should be nearly constant. In addition, we show that in a magnitude-limited survey, if the intrinsic rate of TDEs is independent of black hole mass, the detected events will preferentially have black hole masses $\sim 10^{6.3\pm 0.3}M_{\odot}$ and stellar masses $\sim 1M_{\odot}$, with the width of the mass distribution for disrupted stars sensitive to the stellar mass function in the host galaxy’s center.
@@ 1 Introduction
Tidal disruption events (TDEs) are inherently dramatic: a star is ripped apart by the gravity of a supermassive black hole. The result is a flare in which, for a month or two, the former star radiates with $\sim 10^{10}\times$ its ordinary luminosity. Over the duration of the flare (typically months), TDEs are among the most luminous transients known. Once hard to find, there are now $\sim 100$ known examples (Blanton et al., 2017; Shappee et al., 2014; Kaiser et al., 2002; Bellm et al., 2019; Hammerstein et al., 2023a), and many more can be expected from new instruments (e.g., the Rubin Observatory: Bricman & Gomboc 2020; Szekerczes et al. 2024 and ULTRASAT: Ben-Ami et al. 2022) due to go online soon.
@@
To identify where and how the radiation is emitted, the first task is to track where the debris mass goes. When the star’s orbit is effectively parabolic and the star has been fully disrupted, half the star’s mass is unbound and escapes to infinity, while the other half remains bound to the black hole (Rees, 1988). Although there is a firm consensus that the bound mass is initially placed on highly-eccentric orbits whose apocenters are thousands of gravitational radii, $\sim O(10^{2})\times$ the stellar pericenter, hitherto there has been little agreement about its whereabouts following its first return to the pericenter. The oldest, and still most prevalent, view has been that the bound material immediately forms a compact accretion disk on the pericenter scale Rees (1988). This picture predicts a very luminous X-ray source, but the observed spectra have luminosities $\sim 10^{-2}\times$ this prediction, and shapes well fit by black bodies with temperatures $\gtrsim 10^{4}$ K and radiating areas corresponding to the apocenter scale (see van Velzen et al. (2020); Hammerstein et al. (2023a) for two recent data compilations). To explain these unexpected observations, many (e.g., Strubbe & Quataert (2009, 2011); Metzger & Stone (2016)) have suggested that the X-rays are reprocessed by different material associated with the event whose photosphere happens to be at this distance from the black hole. Alternatively, it has also been suggested (Shiokawa et al., 2015) that, rather than forming a compact disk, the bound material returns to the apocenter region and, upon colliding with later-returning bound debris, forms a large irregular cloud on a scale similar to the apocenter distance. In this picture (Piran et al., 2015), the shocks associated with these collisions power the flare.
@@
Recently, the numerical barriers have been overcome by three different groups (Ryu et al., 2023b; Steinberg & Stone, 2024; Price et al., 2024) using very different numerical methods, but all of them employing well-understood physics. In this paper, we will show that these simulations largely agree on where the bound debris goes after its first return to the vicinity of the black hole: $>99\%$ is placed on orbits extending to thousands of gravitational radii, while $<1\%$ is confined within distances comparable to the disrupted star’s pericenter.
In addition, we will outline the striking observational implications stemming from this consensus. Without the use of any adjustable parameters, they are in good agreement with typical observed values of the luminosity and spectral shape at the peak of the flare, the total radiated energy during the flare peak, the character of the long-term lightcurve, and the stellar and black hole masses most commonly inferred from detected events.
However, it is important to note that this model, although applicable to the majority of the TDE population, does not apply when the black hole mass $M_{\rm BH}\gtrsim 10^{7}M_{\odot}$. For these more massive black holes, the maximum stellar pericenter at which a total disruption takes place is $\lesssim 10r_{g}$ ($r_{g}\equiv GM_{\rm BH}/c^{2}$) (Ryu et al., 2020c). In such cases, relativistic apsidal precession is so strong that it qualitatively changes the characteristic behavior of the debris.
We will first briefly summarize the relevant observational results, focusing on TDEs of the most commonly seen variety, those dominated by an optical/UV flare (Sec. 2). Next (in Sec. 3) we will remind readers of the tidal debris’ characteristic lengthscales, orbital energies, etc. Following these presentations of background, we will summarize the three new global simulations (Sec. 4). The heart of our work will appear in Sec. 5, where we demonstrate how the shared results of the simulations lead to important statements about both the underlying dynamics of TDEs and many of their observed properties. Our conclusions will then be summarized in Sec. 6.
@@ 2.1 The optical flare
In most optical/UV TDEs, the optical/UV flux rises quickly (over a few weeks) to a peak of $\sim 3\times 10^{43}-3\times 10^{44}$ erg s-1, followed by a lengthy (a few months) decline. The spectra are well-fit by single-temperature Planck functions, whose typical temperatures are $\sim 3\times 10^{4}$ K and corresponding black-body emitting radii are $10^{14-15}$ cm (van Velzen et al., 2020; Hammerstein et al., 2023a). When lines are observed, their typical width corresponds to $\sim 5000$ km/sec (see e.g., for a review Gezari, 2021). For a black hole mass $\sim 10^{6}M_{\odot}$, the scale of the emitting area is consistent with the scale of orbits with the speed inferred from the lines.
In some cases, the luminosity declines from its peak $\propto t^{-\alpha}$ with $\alpha\simeq 5/3$, the expected slope of the mass infall rate’s decline (Rees, 1988; Phinney, 1989). However, the entire sample exhibits a wider range of lightcurve power-law indices, $1\lesssim\alpha\lesssim 3$ (Hammerstein et al., 2023a). Approximate integration of lightcurves yields a total radiated energy $\sim 10^{50.5\pm 0.5}$ erg (the ±$\pm 0.5$ in the exponent refers to the standard deviation in the logarithm). At late times, years or more after the peak, the decay is much shallower. For those events observed $\gtrsim 0.5$ yr past the peak, although there is a good deal of scatter, the bolometric luminosity typically falls below the peak by a factor $\sim 10$ and the lightcurve levels out (van Velzen et al., 2019a, 2021; Yao et al., 2023).The late-time radiated energy is therefore comparable to the prompt radiated energy.
@@ 2.2 X-rays
Although there are a few dozen examples of TDEs whose peak X-ray luminosities are comparable to the peak optical/UV luminosities often seen, relatively few TDEs discovered by optical/UV flaring have been associated with X-ray flares, and when there are detectable X-rays during the brightest part of the flare, they are generally much less luminous than the optical/UV (Saxton et al., 2021; Hammerstein et al., 2023a). Nonetheless, the great majority of TDEs seen to produce X-rays, whether optically-discovered or not, have rather soft spectra during the flare: when fitted by a Planck function, the characteristic temperature $kT\sim 50-100$ eV (Saxton et al., 2021). There is, however, a good deal of diversity in their lightcurves: in terms of how rapidly the X-rays decline (Auchettl et al., 2017), non-monotonic behavior, and separation in time from any associated optical/UV flare (Jonker et al., 2020; Malyali et al., 2024; Wevers et al., 2021, 2024) or spectral changes (Jonker et al., 2020; Sazonov et al., 2021; Khorunzhev et al., 2022).
@@ 3.1 Lengthscales
where $m_{\star}$ is the stellar mass in solar masses, $m_{\rm BH,6}$ is the black hole mass in units of $10^{6}M_{\odot}$, and $r_{\rm g}\equiv GM_{\rm BH}/c^{2}$ is the black hole’s gravitational radius. Here and in the rest of the text, we describe the main-sequence mass-radius relation by the power-law $R_{*}=0.93R_{\odot}m_{\star}^{0.88}$ (Ryu et al., 2020d). Whenever $m_{\star}$ appears in a scaling relation, if its exponent is written as a decimal quantity, part of the $m_{\star}$-dependence is through $R_{*}(m_{\star})$.
The tidal radius estimator $r_{\rm t}$ is based on an order-of-magnitude argument. Other radii relevant to tidal disruptions differ from it by factors of order unity. In particular, the critical radius within which a star can be completely disrupted is ${\cal R}_{T}=\Psi(M_{\star},M_{{\rm BH}})r_{\rm t}$ (Ryu et al., 2020d, e). The correction factor $\Psi(M_{\star},M_{{\rm BH}})$ is defined in the Appendix. Although $\Psi(M_{\star},M_{{\rm BH}})$ is of order unity, its dependence on the stellar mass and the black hole mass is important when estimating TDE rates.
@@ 3.2 Energy scales and orbital properties
In other words, the orbital energy of the debris is, for typical parameters, $\sim 10^{-2}$ the potential energy near the nominal tidal radius. However, once again, detailed calculations (Ryu et al., 2020d, e; Law-Smith et al., 2020) have found that consideration of the internal density profile of main sequence stars introduces order-unity correction factors: Ryu et al. (2020d) defined them by $\Xi=\Delta E/\Delta E_{0}$, where the energy range $-\Delta E<E<+\Delta E$ contains 90% of the debris mass. Details about the correction factor $\Xi$, including fitting formulæ for its dependence on $M_{\star}$ and $M_{{}_{\rm BH}}$, are provided in the Appendix  A.
@@
As $m_{\star}$ increases, the internal density profile of main sequence stars becomes increasingly centrally-concentrated. This causes $\Xi$ to rise as a function of $M_{\star}$, and the net result is for there to be almost no net trend in $t_{0}$ as $M_{\star}$ increases from $\simeq 0.1M_{\odot}$ to $\simeq 10M_{\odot}$ (Ryu et al., 2020d). On the other hand, the explicit scaling with $M_{\rm BH}$ is augmented by the implicit dependence on $M_{\rm BH}$ through $\Xi$, making $t_{0}\propto m_{{}_{\rm BH,6}}^{0.6}$.
@@ 3.3 Circularization and the inverse energy crisis
where $\sigma_{\rm SB}$ is the Stefan-Boltzmann constant. This is 50$\sim$ 50$\times$ 50 × the observed optical/UV temperature (van Velzen et al., 2020; Hammerstein et al., 2023a). Typical observed velocities from motion at a few tens of $r_{\rm g}$ would be $\sim 30,000$ km s-1, about ten times larger than implied by the measured emission line widths.
@@ 4.1 Realistic, Global Simulations
Three simulations stand out as satisfying all these criteria. We list them in order of their publication dates. Ryu et al. (2023b) employed two fixed grids, one a small box following the star’s orbit, the other a large volume surrounding the black hole. Simultaneous evolution on the two grids was coordinated through a “multipatch" system (Shiokawa et al., 2018; Avara et al., 2024). The programs running on both grids solve the general relativistic hydrodynamics equations; the one responsible for the star adds relativistically-consistent stellar self-gravity. Ryu et al. (2023b) considered a $3M_{\odot}$ star, whose internal structure was taken from a MESA model of a middle-aged main-sequence star, and a $10^{5}M_{\odot}$ black hole. The equation of state included LTE radiation pressure, but there was no computation of radiation transfer because, for these parameters, the cooling time was always very long compared to the evolution time. The duration of this simulation was $3t_{0}$. Steinberg & Stone (2024) solved the equations of Newtonian hydrodynamics and radiation transport in the flux-limited diffusion approximation on a moving mesh. Stellar self-gravity was computed via a quadrupole moment tree, and the black hole’s gravity was approximated as a Paczynski-Wiita potential with a softening term. They chose a $1M_{\odot}$ star and a $10^{6}M_{\odot}$ black hole, but described the internal structure of the star as an $n=3/2$ polytrope, i.e., an isentropic structure for a gas with an adiabatic index of 5/3. This simulation ran for $\approx 1.4t_{0}$. The third (Price et al., 2024) was produced by an SPH code with relativistic hydrodynamics in a Schwarzschild spacetime and the same equation of state as in Ryu et al. (2023b). Their star’s initial structure was the same as in Steinberg & Stone (2024), but its self-gravity was calculated as an integral over the Newtonian Green’s function for the Poisson Equation with a softening term to smooth short lengthscale fluctuations. The simulation covered a time up to $\simeq 9t_{0}$. All three of these simulations provided resolution capable of describing the smallest relevant structures, whether in the disrupting star or in shocks within the debris flow. Steinberg & Stone (2024) presented an especially detailed discussion of these issues, but the strong agreement in results between the three suggests that the other two were comparably well-resolved.
@@ 4.2 Consensus dynamical results
Debris that has passed through the nozzle shock swings back out toward an apocenter that is slightly smaller because of the orbital energy dissipated in the shock; its natal specific angular momentum is so small that the angular momentum exchange in the shock hardly affects the apocenter. Because the disruption of the star takes place across a range of distances from the black hole (roughly from $r\simeq r_{\rm p}$ to $r\simeq 20r_{\rm p}$: Ryu et al. (2020e)) and because of relativistic effects in both the tidal stress and the orbits, the lines of apsides of the debris orbits stretch across a range of angles $\sim 10^{\circ}$ (Shiokawa et al., 2015). When the range of angles is this small and the orbits are highly eccentric, the streams intersect near apocenter (see also the related geometric argument of Dai et al. (2015)). Two views of these shocks are presented in Figure 1, which highlights shock locations by portraying temperature (see also Fig. 1 of Steinberg & Stone 2024).